# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v2

@@ Abstract
Temporally causal representation learning aims to identify the latent causal process from time series observations, but most methods require the assumption that the latent causal processes do not have instantaneous relations. Although some recent methods achieve identifiability in the instantaneous causality case, they require either interventions on the latent variables or grouping of the observations, which are in general difficult to obtain in real-world scenarios. To fill this gap, we propose an IDentification framework for instantaneOus Latent dynamics (IDOL) by imposing a sparse influence constraint that the latent causal processes have sparse time-delayed and instantaneous relations. Specifically, we establish identifiability results of the latent causal process based on sufficient variability and the sparse influence constraint by employing contextual information of time series data. Based on these theories, we incorporate a temporally variational inference architecture to estimate the latent variables and a gradient-based sparsity regularization to identify the latent causal process. Experimental results on simulation datasets illustrate that our method can identify the latent causal process. Furthermore, evaluations on multiple human motion forecasting benchmarks with instantaneous dependencies indicate the effectiveness of our method in real-world settings.
@@ 1 Introduction
Time series analysis [1, 2, 3, 4, 5], which has been found widespread applications across diverse fields such as weather [6, 7], finance [8, 9], and human activity recognition [10, 11], aims to capture the underlying patterns [12, 13] behind the time series data. To achieve this goal, one solution is to estimate the latent causal processes [14, 15, 16]. However, without further assumptions [17, 18], it is challenging to identify latent causal processes, i.e., ensure that the estimated latent causal process is correct.
@@
Researchers have exploited Independent Component Analysis (ICA) [19, 20, 21, 22], where observations are generated from latent variables via a linear mixture function, to identify the latent causal process. To deal nonlinear cases, different types of assumptions including sufficient changes [23, 24, 25] and structural sparsity [26, 27] are proposed to meet the independent change of latent variables. Specifically, several works leverage auxiliary variables to achieve strong identifiable results of latent variables [28, 29, 18, 30, 24, 31]. To seek identifiability in an unsupervised manner, other researchers [32, 33] propose the structural sparsity on the generation process from latent sources to observation. However, these methods usually do not have instantaneous relations, e.g., the latent variables are mutually independent conditional on their time-delayed parents at previous timestamps. Hence, further assumptions are used for identifiability. For example, Lippe et al. [34] demonstrate identifiability by assuming that there exist interventions on latent variables, and Morioka et al. [35] yield identifiability with the assumption of the grouping of observational variables. Recently, Zhang et al. [36] use the sparse structures of latent variables to achieve identifiability of static data under distribution shift. Beyond it, in this work, we demonstrate that temporal identifiability benefits from the sparsity constraint on the causal influences in the latent causal processes by employing the temporal contextual information, i.e., the variables across multiple timestamps. Please refer to Appendix A1 for further discussion of related works, including the identification of latent variables and instantaneous time series.
@@
Figure 1: Three different data generation processes with time-delayed and instantaneous dependencies. (a) iCITRIS [34] requires intervention variables $\textbf{i}_{t}$ for latent variables (the gray nodes with blue rim.). (b) G-CaRL [35] assumes that observed variables can be grouped according to which latent variables they are connected to (the blue dotted lines), (c) IDOL requires a sparse latent causal process (the blue solid lines), which is close to real-world scenarios.
@@
Although these methods [34, 35] have demonstrated the identifiability of latent variables with instantaneous dependencies, they often impose strict assumptions on the latent variables and observations. Specifically, as shown in Figure 1 (a), iCITRIS [34] reaches identifiability by assuming that interventions $\bm{I}_{t}$, i.e., the gray nodes with blue rim, on the latent variables, which is expensive and difficult in practice. For instance, considering the joints as the latent variables of a human motion process, it is almost impossible to intervene in the joints of the knee when a human is walking. Moreover, as shown in Figure 1 (b), G-CaRL [35] employs the grouping of observations to identify the latent variables, i.e., the blue dotted lines. However, the grouping of observations is usually expensive, for example, grouping the sensors on a human body from the joints of skeletons requires massive of expert knowledge and labor. Therefore, it is urgent to find a more general approach with an appropriate assumption to identify the latent causal process with instantaneous dependencies.
@@
To identify the latent variables in a more relaxed assumption that is applicable to real-world time series data, we present a general IDentification framework for instantaneOus Latent dynamics (IDOL) with the only assumption of the sparse causal influences in the latent processed named “sparse latent process” as shown in Figure 1 (c), where the transition process denoted by the blue solid lines is sparse. It is worth noting that the sparse latent process assumption is common and naturally appears in many practical time series data. In the example of human motion forecasting, a skeleton of the human body, which can be considered to be a latent causal process, is usually sparse, for example, the changes of arms and legs can be treated as independent. Based on the latent process sparsity, we first establish the relationship between the ground truth and the estimated latent variables by employing sufficient variability. Sequentially, we propose the sparse latent process assumption to achieve the component-wise identifiability of latent variables by leveraging the contextual information of time series data. Building upon this identifiability framework, we develop the IDOL model, which involves a temporally variational inference neural architecture to reconstruct the latent variables and a gradient-based sparsity penalty to encourage the sparse causal influences in the latent processe for identifiability. Evaluation results on the simulation datasets support the theoretical claims and experiment results on several benchmark datasets for time series forecasting highlight the effectiveness of our method.
@@ 2.1 Time Series Generative Model under Instantaneous Latent Causal Process
We begin with the data generation process as shown in Figure 1 (c), where the latent causal process is sparse. Suppose that we have time series data with discrete timestamps $\textbf{X}=\{{\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t},\cdots,{\mathbf{x}}_{T}\}$, where ${\mathbf{x}}_{t}$ is generated from latent variables ${\mathbf{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$ via an invertible and nonlinear mixing function ${\mathbf{g}}$, which is shown in Equation 1,
$\small{\mathbf{x}}_{t}={\mathbf{g}}({\mathbf{z}}_{t}).$ (1)
Specifically, the $i$-th dimension of the latent variable $z_{t,i}$ is generated via the latent causal process, which is assumed to be related to the time-delayed parent variables $\text{Pa}_{d}(z_{t,i})$ and the instantaneous parents $\text{Pa}_{e}(z_{t,i})$, respectively. Formally, it can be written via a structural equation model [37] as
$\small\begin{split}z_{t,i}=f_{i}(\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}),\epsilon_{t,i})\quad\text{with}\quad\epsilon_{t,i}\sim p_{\epsilon_{i}},\end{split}$ (2)
where $\epsilon_{t,i}$ denotes the temporally and spatially independent noise extracted from a distribution $p_{\epsilon_{i}}$. To better understand this data generation process, we take an example of human motion forecasting, where the joint positions can be considered as latent variables. In this case, the rigid body effects among joints result in instantaneous effects, and the kinematic representations like relative joint angles between limbs recorded from the sensors on the human body are observed variables that are the mixtures of joint positions. Therefore, we can consider the process from latent joint positions to the observed kinematic representations as the mixing process.
@@ 2.1.1 Identifiability of Latent Causal Processes
Based on the aforementioned generation process, we further provide the definition of the identifiability of latent causal process with instantaneous dependency in Definition 1. Moreover, if the estimated latent processes can be identified at least up to permutation and component-wise invertible transformation, the latent causal relations are also immediately identifiable because conditional independence relations fully characterize instantaneous causal relations in an instantaneous causally sufficient system, where there are no causal confounders in the latent causal processes.
Definition 1 (Identifiable Latent Causal Process).
Let $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}$ be a sequence of observed variables generated by the true latent causal processes specified by $(f_{i},p({\epsilon_{i}}),\mathbf{g})$ given in Eq 1 and 2. A learned generative model $(\hat{f}_{i},\hat{p}({\epsilon_{i}}),\hat{\mathbf{g}})$ is observational equivalent to $(f_{i},p({\epsilon_{i}}),\mathbf{g})$ if the model distribution $p_{\hat{f}_{i},\hat{p}_{\epsilon},\hat{\mathbf{g}}}(\{\mathbf{x}\}_{t=1}^{T})$ matches the data distribution $p_{f_{i},p_{\epsilon},\mathbf{g}}(\{\mathbf{x}\}_{t=1}^{T})$ for any value of $\mathbf{x}_{t}$. We say latent causal processes are identifiable if observational equivalence can lead to a version of generative process up to a permutation $\pi$ and component-wise invertible transformation $\mathcal{T}$:
$\small p_{\hat{f}_{i},\hat{p}_{\epsilon_{i}},\hat{\mathbf{g}}}\big{(}\{\mathbf{x}\}_{t=1}^{T}\big{)}=p_{f_{i},p_{\epsilon_{i}},\mathbf{g}}\big{(}\{\mathbf{x}\}_{t=1}^{T}\big{)}\ \ \Rightarrow\ \ \hat{\mathbf{g}}=\mathbf{g}\circ\pi\circ\mathcal{T}.$ (3)
@@
$\small\mathbf{\hat{z}}_{t}=\mathbf{\hat{g}}^{-1}(\mathbf{x}_{t})=\big{(}\mathcal{T}^{-1}\circ\pi^{-1}\circ\mathbf{g}^{-1}\big{)}(\mathbf{x}_{t})=\big{(}\mathcal{T}^{-1}\circ\pi^{-1}\big{)}\big{(}\mathbf{g}^{-1}(\mathbf{x}_{t})\big{)}=\big{(}\mathcal{T}^{-1}\circ\pi^{-1}\big{)}(\mathbf{z}_{t}).$ (4)
Furthermore, with identified latent variables, the whole causal process within latent variables can also be determined since intervention can be applied to any single variable easily.
@@ 3 Identification Results for Latent Causal Process
Given the definition of identification of latent causal process, we show how to identify the latent causal process in Figure 1 (c) under a sparse latent process. Specifically, we first leverage the connection between conditional independence and cross derivatives [38] and the sufficient variability of temporal data to discover the relationships between the estimated and the true latent variables, which is shown in Theorem 1. Moreover, we establish the identifiability result of latent variables by enforcing the sparse causal influences, as shown in Theorem 2. Finally, we also show that the existing identifiability results of the temporally causal representation learning methods are a special case of our IDOL method, which is shown in Corollary 2.1.
@@ 3.1 Relationships between Ground-truth and Estimated Latent Variables
In this section, we figure out the relationships between ground truth and estimated latent variables under a temporal scenario with instantaneous dependence. In order to incorporate contextual information to aid in the identification of latent variables ${\mathbf{z}}_{t}$, latent variables of $L$ consecutive timestamps including ${\mathbf{z}}_{t}$, are taken into consideration. Without loss of generality, we consider a simplified case, where the length of the sequence is 2, i.e., $L=2$, and the time lag is 1, i.e., $\tau=1$. Please refer to Appendix A2.2 for the general case of multiple lags and sequence lengths
@@
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ with the corresponding process $\hat{f}_{i},\hat{p}(\epsilon),\hat{g}$, where $\hat{g}$ is invertible, suppose the process subject to observational equivalence $\mathbf{x}_{t}=\mathbf{\hat{g}}(\mathbf{\hat{z}}_{t})$. Let $\mathbf{c}_{t}\triangleq\{\mathbf{z}_{t-1},\mathbf{z}_{t}\}$ and $\mathcal{M}_{\mathbf{c}_{t}}$ be the variable set of two consecutive timestamps and the corresponding Markov network respectively. Suppose the following assumptions hold:
  • A1 (Smooth and Positive Density): The probability function of the latent variables ${\mathbf{c}}_{t}$ is smooth and positive, i.e., $p_{{\mathbf{c}}_{t}}$ is third-order differentiable and $p_{{\mathbf{c}}_{t}}>0$ over $\mathbb{R}^{2n}$,
@@
    $\small\begin{split}w(m)=&\Big{(}\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}^{2}\partial z_{t-2,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}^{2}\partial z_{t-2,m}}\Big{)}\oplus\\ &\Big{(}\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}\partial z_{t-2,m}},\cdots,\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}\partial z_{t-2,m}}\Big{)}\oplus\Big{(}\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}\partial z_{t-2,m}}\Big{)}_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})},\end{split}$ (5)
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})$ denotes all pairwise indice such that $c_{t,i},c_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $4n+2|\mathcal{M}_{{\mathbf{c}}_{t}}|$ different values of ${\mathbf{z}}_{t-2,m}$, as the $4n+2|\mathcal{M}_{{\mathbf{c}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
Then for any two different entries $\hat{c}_{t,k},\hat{c}_{t,l}\in\hat{{\mathbf{c}}}_{t}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ over estimated $\hat{{\mathbf{c}}}_{t}$,
(i) Each ground-truth latent variable $c_{t,i}\in{\mathbf{c}}_{t}$ is a function of at most one of $\hat{c}_{k}$ and $\hat{c}_{l}$,
(ii) For each pair of ground-truth latent variables $c_{t,i}$ and $c_{t,j}$ that are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$ over ${\mathbf{c}}_{t}$, they can not be a function of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$ respectively.
Proof Sketch. The proof can be found in Appendix A2.1. First, we establish a bijective transformation between the ground-truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$ to connect them together. Next, we utilize the structural properties of the ground-truth Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$ to constrain the structure of the estimated Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ through the connection between conditional independence and cross derivatives[38], i.e., $c_{t,i}\perp c_{t,j}|{\mathbf{c}}_{t}\textbackslash\{c_{t,i},c_{t,j}\}$ implies $\frac{\partial^{2}\log p({\mathbf{c}}_{t})}{\partial c_{t,i}\partial c_{t,j}}=0$. By introducing the sufficient variability assumption, we further construct a linear system with a full rank coefficient matrix to ensure that the only solution holds, i.e.,
$\small\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}=0,\quad\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}=0,\quad\frac{\partial^{2}c_{t,i}}{\partial c_{t,k}\partial c_{t,l}}=0,$ (6)
@@
Discussion. The first assumptions from Theorem 1 are standard in the component-wise identification of existing nonlinear ICA for time series data[39], i.e., the latent variables are changing continuously, which is common in time series data. The second assumption has been commonly adopted for the existing identifiable results of temporally causal representation learning for sufficient changes. Theorem 1 provides an insight, such that when observational equivalence holds, the variables that are directly related in the true Markov network must be functions of directly related variables in the estimated Markov network. Please note that $\tau$ and $L$ can be easily generalized to any value by making some modifications on the assumption $A2$. The $\tau$ timestamps that are conditioned on provide sufficient changes, and $L$ sequence length provides a sparse structure, which will be discussed in Theorem 2. A detailed discussion is given in Appendix A2.2.

### v3

@@ Abstract
Temporally causal representation learning aims to identify the latent causal process from time series observations, but most methods require the assumption that the latent causal processes do not have instantaneous relations. Although some recent methods achieve identifiability in the instantaneous causality case, they require either interventions on the latent variables or grouping of the observations, which are in general difficult to obtain in real-world scenarios. To fill this gap, we propose an IDentification framework for instantaneOus Latent dynamics (IDOL) by imposing a sparse influence constraint that the latent causal processes have sparse time-delayed and instantaneous relations. Specifically, we establish identifiability results of the latent causal process up to a Markov equivalence class based on sufficient variability and the sparse influence constraint by employing contextual information. We further explore under what conditions the identification can be extended to the causal graph. Based on these theoretical results, we incorporate a temporally variational inference architecture to estimate the latent variables and a gradient-based sparsity regularization to identify the latent causal process. Experimental results on simulation datasets illustrate that our method can identify the latent causal process. Furthermore, evaluations on human motion forecasting benchmarks indicate the effectiveness in real-world settings. Source code is available at https://github.com/DMIRLAB-Group/IDOL.
@@ 1 Introduction
Time series analysis (zhang2023self; tang2021probabilistic; li2023difformer; wu2022timesnet; luo2024moderntcn), which has been found widespread applications across diverse fields such as weather (bi2023accurate; wu2023interpretable), finance (tsay2005analysis; huynh2022few), and human activity recognition (yang2015deep; kong2022human), aims to capture the underlying patterns (wang2019time; jin2022multivariate) behind the time series data. To achieve this, one solution is to estimate the latent causal processes (tank2021neural; li2023transferable; zheng2019using). However, without further assumptions (locatello2020sober), it is challenging to identify latent causal processes, i.e., ensure that the estimated latent causal process is correct.
@@
Researchers have exploited Independent Component Analysis (ICA) (hyvarinen2000independent; comon1994independent; hyvarinen2013independent; zhang2007kernel), where observations are generated from latent variables via a linear mixing function, to identify the latent causal process. To deal with nonlinear cases, different types of assumptions including sufficient changes (yao2021learning; khemakhem2020variational; xie2022multi) and structural sparsity (lachapelle2022disentanglement; zheng2022identifiability) are proposed to meet the independent change of latent variables. Specifically, several works leverage auxiliary variables to achieve strong identifiable results of latent variables (halva2020hidden; halva2021disentangling; hyvarinen2016unsupervised; hyvarinen2017nonlinear; khemakhem2020variational; hyvarinen2019nonlinear). To seek identifiability in an unsupervised manner, other researchers (ng2024identifiability; zheng2024generalizing) propose the structural sparsity on the generation process from latent sources to observation. However, these methods usually do not have instantaneous relations, e.g., the latent variables are mutually independent conditional on their time-delayed parents at previous timestamps. Hence, further assumptions are used for identifiability. For example, lippe2023causal demonstrate identifiability by assuming that there exist interventions on latent variables, and morioka2023causal yield identifiability with the assumption of the grouping of observational variables. Recently, zhang2024causal use the sparse structures of latent variables to achieve identifiability of static data under distribution shift. Beyond it, in this work, we demonstrate that temporal identifiability benefits from the sparsity constraint on the causal influences in the latent causal processes by employing the temporal contextual information, i.e., the variables across multiple timestamps. Please refer to Appendix A and D for further discussion of related works, including the identification of latent variables and instantaneous time series.
@@
Figure 1: Three different data generation processes with time-delayed and instantaneous dependencies. (a) iCITRIS (lippe2023causal) requires intervention variables $\textbf{i}_{t}$ for latent variables (the gray nodes with blue rim.). (b) G-CaRL (morioka2023causal) assumes that observed variables can be grouped according to which latent variables they are connected to (the blue dotted lines), (c) IDOL requires a sparse latent causal process (the blue solid lines).
@@
Although these methods (lippe2023causal; morioka2023causal) have demonstrated the identifiability of latent variables with instantaneous dependencies, they often impose strict assumptions on the latent variables and observations. Specifically, as shown in Figure 1 (a), iCITRIS (lippe2023causal) reaches identifiability by assuming that interventions $\bm{I}_{t}$, i.e., the gray nodes with blue rim, on the latent variables, which is expensive and difficult in practice. For instance, considering the joints as the latent variables of a human motion process, it is almost impossible to intervene in the joints of the knee when a human is walking. Moreover, as shown in Figure 1 (b), G-CaRL (morioka2023causal) employs the grouping of observations to identify the latent variables, i.e., the blue dotted lines. However, the grouping of observations is usually expensive, for example, grouping the sensors on a human body from the joints of skeletons requires massive of expert knowledge and labor. Therefore, it is urgent to find a more general approach with an appropriate assumption to identify the latent causal process with instantaneous dependencies.
@@
To identify the latent variables in more relaxed assumptions, we present a general IDentification framework for instantaneOus Latent dynamics (IDOL) with the only assumption of the sparse causal influences in the latent processed named “sparse latent process” as shown in Figure 1 (c), where the transition process denoted by the blue solid lines is sparse. It is worth noting that the sparse latent process assumption is common and naturally appears in many practical time series data. In the example of human motion forecasting, a skeleton of the human body, which can be considered to be a latent causal process, is usually sparse. This is because there are a few connections among the joints of the human body. Based on the latent process sparsity, we first establish the relationship between the ground truth and the estimated latent variables by employing sufficient variability. Sequentially, we propose the sparse latent process assumption to achieve the component-wise identifiability of latent variables by leveraging the contextual information of time series data, which implies the identification of Markov equivalence class. Furthermore, we can extend to the identification of the causal graph when the endpoints of instantaneous edges do not share identical time-delayed parents, i,e, the example in Figure 1 (c) precisely satisfies this condition. Building upon this identifiability framework, we develop the IDOL model, which involves a variational inference neural architecture to reconstruct the latent variables and a gradient-based sparsity penalty to encourage the sparse causal influences in the latent process for identifiability. Evaluation results on the simulation datasets support the theoretical claims and experiment results on several benchmark datasets for time series forecasting highlight the effectiveness of our method.
@@ 2.1 Time Series Generative Model under Instantaneous Latent Causal Process
We begin with the data generation process as shown in Figure 1 (c) where the latent causal process is sparse. To facilitate clarity, we adopt the terminology widely used in ICA literature like observation ${\mathbf{x}}_{t}$, latent variables ${\mathbf{z}}_{t}$ and mixing function $g$ (hyvarinen2016unsupervised). Suppose that we have time series data with discrete timestamps $\textbf{X}=\{{\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t},\cdots,{\mathbf{x}}_{T}\}$, where ${\mathbf{x}}_{t}$ is generated from latent variables ${\mathbf{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$ via an invertible and nonlinear mixing function $\mathbf{g}$}. We have:
${\mathbf{x}}_{t}={\mathbf{g}}({\mathbf{z}}_{t}).$ (1)
Specifically, the $i$-th dimension of the latent variable $z_{t,i}$ is generated via the latent causal process, which is assumed to be related to the time-delayed parent variables $\text{Pa}_{d}(z_{t,i})$ and the instantaneous parents $\text{Pa}_{e}(z_{t,i})$. Formally, it can be written via a structural equation model (neuberg2003causality) as
$\begin{split}z_{t,i}=f_{i}(\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}),\epsilon_{t,i})\quad\text{with}\quad\epsilon_{t,i}\sim p_{\epsilon_{i}},\end{split}$ (2)
where $\epsilon_{t,i}$ denotes the temporally and spatially independent noise extracted from a distribution $p_{\epsilon_{i}}$. To better understand this data generation process, we take an example of human motion forecasting, where the joint positions can be considered as latent variables. In this case, the rigid body effects among joints result in instantaneous effects, and the human movement trajectory recorded by the sensor are observed variables. As a result, we can consider the process from latent joint positions to movement trajectory as the mixing process.
@@ 2.1.1 Identifiability of Latent Causal Processes
Based on the aforementioned generation process, we further provide the definition of the identifiability of latent causal process with instantaneous dependency in Definition 1. Moreover, if the estimated latent processes can be identified at least up to permutation and component-wise invertible transformation, the latent causal relations are also immediately identifiable up to a Markov equivalence class (spirtes2001causation). We further show how to go beyond the Markov equivalence class and identify the instantaneous causal relations with a mild assumption in Theorem 3.
Definition 1 (Identifiable Latent Causal Process (yao2022temporally; yao2021learning)).
[…]
$p_{\hat{f}_{i},\hat{p}_{\epsilon_{i}},\hat{\mathbf{g}}}\big(\{\mathbf{x}\}_{t=1}^{T}\big)=p_{f_{i},p_{\epsilon_{i}},\mathbf{g}}\big(\{\mathbf{x}\}_{t=1}^{T}\big)\ \ \Rightarrow\ \ \hat{\mathbf{g}}=\mathbf{g}\circ\pi\circ\mathcal{T}.$ (3)
@@
$\mathbf{\hat{z}}_{t}=\mathbf{\hat{g}}^{-1}(\mathbf{x}_{t})=\big(\mathcal{T}^{-1}\circ\pi^{-1}\circ\mathbf{g}^{-1}\big)(\mathbf{x}_{t})=\big(\mathcal{T}^{-1}\circ\pi^{-1}\big)\big(\mathbf{g}^{-1}(\mathbf{x}_{t})\big)=\big(\mathcal{T}^{-1}\circ\pi^{-1}\big)(\mathbf{z}_{t}).$ (4)
@@ 3 Identification Results for Latent Causal Process
Given the definition of identification of latent causal process, we show how to identify the latent causal process in Figure 1 (c) under a sparse latent process. Please note that our theorem is discussed under Markov Network. For more details, please refer to Appendix B.6. Specifically, we first leverage the connection between conditional independence and cross derivatives (lin1997factorizing) and the sufficient variability of temporal data to discover the relationships between the estimated and the true latent variables, which is shown in Theorem 1. Moreover, we establish the identifiability result of latent variables by enforcing the sparse causal influences, as shown in Theorem 2. Moreover, we also show that the existing identifiability results of the temporally causal representation learning methods are a special case of our IDOL method, which is shown in Corollary 1. Finally, under a mild assumption, we achieve identification of the causal graph underlying the latent causal process, as shown in Theorem 3.
@@ 3.1 Relationships between Ground-truth and Estimated Latent Variables
In this section, we figure out the relationships between ground truth and estimated latent variables under a temporal scenario with instantaneous dependence. In order to incorporate contextual information to aid in the identification of latent variables ${\mathbf{z}}_{t}$, latent variables of $L$ consecutive timestamps including ${\mathbf{z}}_{t}$, are taken into consideration. Without loss of generality, we consider a simplified case, where the length of the sequence is 2, i.e., $L=2$, and the time lag is 1, i.e., $\tau=1$. Please refer to Appendix B.2 for the general case of multiple lags and sequence lengths.
@@
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ with the corresponding process $\hat{f}_{i},\hat{p}(\epsilon),\hat{g}$, where $\hat{g}$ is invertible, suppose that the process subject to observational equivalence $\mathbf{x}_{t}=\mathbf{\hat{g}}(\mathbf{\hat{z}}_{t})$. Let $\mathbf{c}_{t}\triangleq\{\mathbf{z}_{t-1},\mathbf{z}_{t}\}\in\mathbb{R}^{2n}$ and that $\mathcal{M}_{\mathbf{c}_{t}}$ be the variable set of two consecutive timestamps and the corresponding Markov network respectively. Suppose that the following assumptions hold:
  • A1 (Smooth and Positive Density): The conditional probability function of the latent variables ${\mathbf{c}}_{t}$ is smooth and positive, i.e., $p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})$ is third-order differentiable and $p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})>0$ over $\mathbb{R}^{2n}$,
@@
    $\begin{split}w(m)=&\Big(\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}^{2}\partial z_{t-2,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}^{2}\partial z_{t-2,m}}\Big)\oplus\\ &\Big(\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}\partial z_{t-2,m}},\cdots,\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}\partial z_{t-2,m}}\Big)\oplus\Big(\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}\partial z_{t-2,m}}\Big)_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})},\end{split}$ (5)
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})$ denotes all pairwise indice such that $c_{t,i},c_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $2n+|\mathcal{M}_{{\mathbf{c}}_{t}}|$ different values of ${\mathbf{z}}_{t-2,m}$, such that the $2n+|\mathcal{M}_{{\mathbf{c}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
Then for any two different entries $\hat{c}_{t,k},\hat{c}_{t,l}$ of $\hat{{\mathbf{c}}}_{t}\in\mathbb{R}^{2n}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ over estimated $\hat{{\mathbf{c}}}_{t}$,
(i) Each ground-truth latent variable $c_{t,i}$ of ${\mathbf{c}}_{t}\in\mathbb{R}^{2n}$ is a function of at most one of $\hat{c}_{k}$ and $\hat{c}_{l}$,
(ii) For each pair of ground-truth latent variables $c_{t,i}$ and $c_{t,j}$ of ${\mathbf{c}}_{t}\in\mathbb{R}^{2n}$ that are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$ over ${\mathbf{c}}_{t}$, they can not be a function of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$ respectively.
Proof Sketch. The proof can be found in Appendix B.1. First, we establish a bijective transformation between the ground-truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$ to connect them together. Next, we utilize the structural properties of the ground-truth Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$ to constrain the structure of the estimated Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ through the connection between conditional independence and cross derivatives (lin1997factorizing), i.e., $c_{t,i}\perp c_{t,j}|{\mathbf{c}}_{t}\backslash\{c_{t,i},c_{t,j}\}$ implies $\frac{\partial^{2}\log p({\mathbf{c}}_{t})}{\partial c_{t,i}\partial c_{t,j}}=0$. By introducing the sufficient variability assumption, we further construct a linear system with a full rank coefficient matrix to ensure that the only solution holds, i.e.,
$\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}=0,\quad\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}=0,\quad\frac{\partial^{2}c_{t,i}}{\partial c_{t,k}\partial c_{t,l}}=0,$ (6)
@@
Theorem 1 provides an insight, such that when observational equivalence holds, the variables that are directly related in the true Markov network must be functions of directly related variables in the estimated Markov network. Please note that $\tau$ and $L$ can be easily generalized to any value by making some modifications on the assumption $A2$. The $\tau$ timestamps that are conditioned on provide sufficient changes, and $L$ sequence length provides a sparse structure, which will be discussed in Theorem 2. A detailed discussion is given in Appendix B.2.